OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f−1(3)?

OpenStudy (anonymous):

NOTE: The 2 is small towards the bottom, next to log. The -1 is small and positioned like an exponent after f.

OpenStudy (hoblos):

let y=f(x) then we have y=log2 (x + 4) to find y-1 we replace x by y and y by x ---> x=log2 (y + 4) and solve for y do you know how to solve for y here ?

OpenStudy (anonymous):

to multiply 2 and 4?

OpenStudy (ipwnbunnies):

\[\log _{a}b = x \rightarrow 2^{x} = b\]

OpenStudy (hoblos):

\[a ^{x}=b\] @iPwnBunnies

OpenStudy (ipwnbunnies):

^^ Right. I didn't mean to giveaway the answer. D:

OpenStudy (ipwnbunnies):

\[\log _{a}b = x \rightarrow a^{x} = b\]

OpenStudy (hoblos):

lol @iPwnBunnies @mathhelpnow can you use this identity here ?

OpenStudy (anonymous):

I'll try, 2 is a, and 4 is b, right?

OpenStudy (ipwnbunnies):

(y+4) would be b.

OpenStudy (hoblos):